# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 3.4 Weighted regression estimator
A third regression-based augmented estimator is the weighted regression estimator (WET)[31, 26, 32, 33] that, in our case, utilizes balancing weights during outcome regression process. The weighted regression estimator can directly incorporate both propensity scores and survey weights. It enables effective bias correction and aligns the estimated coefficients more closely with the true population parameters. [26, 30] These properties are especially advantageous in complex data environments where both models are prone to misspecification and the sampling process is not necessarily independent of exposure and treatment assignment. Adapted to our setting of survey sampling, the weighted regression estimator can be expressed as,
@@
$\widehat{\tau}_{h}^{\text{WET}}=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})\left(m_{1}^{\text{WET}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{WET}}\right)-m_{0}^{\text{WET}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{WET}}\right)\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})}$ (18)
where $m_{1}^{\text{WET}}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{1}^{\text{WET}}\right)$ and $m_{0}^{\text{WET}}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{0}^{\text{WET}}\right)$ are the outcome regression models with canonical link functions, similar to those in the clever covariate estimator, but using weighted regression with balancing weights $\omega_{1,p}\left(\mathbf{X}\right)$ and $\omega_{0,p}\left(\mathbf{X}\right)$. In this estimator, survey weights are incorporated in three distinct stages: during the survey-weighted propensity score regression for $e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$, for the construction of the balancing weight pairs $\omega_{1,p}(\mathbf{X}_{i})$ and $\omega_{0,p}(\mathbf{X}_{i})$, which subsequently serve as regression weights in the outcome modeling process, and in standardizing the predicted potential outcomes toward the target population. This strategy adjusts the weighted least squares coefficients $\widehat{\boldsymbol{\alpha}}_{1}^{\text{WET}}$ and $\widehat{\boldsymbol{\alpha}}_{0}^{\text{WET}}$ by accounting for the mechanism between $\mathbf{X}$ and potential outcomes under treatment and control conditions respectively in the target population, as long as the SUTVA and weak exchangeability assumptions in Section 2.1 hold. For ease of reference, Table 2 summarizes where survey weights need to be applied for each simple weighting estimator and augmented weighting estimator.
Similar to the moment and clever covariate estimators by sharing the same influence function, the weighted regression estimator ensures consistent estimation when the survey-weighted propensity score model $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$ is correctly specified, with additional robustness depending on the definition of the tilting function. When the tilting function $h(\mathbf{X})\propto 1$, the weighted regression estimator is doubly robust, leading to consistent estimation if either the weighted outcome regression model or the survey-weighted propensity score model $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$ is correctly specified. In cases where $h(\mathbf{X})$ explicitly depends on $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$, incorporating the weighted outcome regression models can enhance efficiency relative to the simple weighting estimators discussed in Section 2. Consistency to the originally intended target estimand in this case is achieved only if $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$ is correctly specified; otherwise, the estimator will converge to a biased target estimand $\widetilde{\tau}_{h}$, determined by the misspecified tilting function $h(e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}}))$.
@@
Compared to the clever covariate estimator, the weighted regression estimator may offer better performance in scenarios involving highly variable survey weights or extreme propensity scores. As discussed by Kang et al. [30]and Robins et al., [26] the weighted regression estimator is more stable under extreme balancing weights, reducing the risk of overfitting that may occur under the clever covariate construction. However, under the IPW scheme, the weighted regression estimator remains susceptible to the impact of extreme inverse propensity scores, which can lead to inflated variance and finite-sample bias, particularly when these extreme weights dominate the regression model. We plan to examine to what extent these observations carry forward to the survey observational study settings in Section 4.
@@
Estimator Estimating propensity scores Constructing the balancing weights Fitting the outcome regression model Standardizing predicted potential outcomes to the target estimand
PSW $\sqrt{}$ $\sqrt{}$ Not applicable Not applicable
MOM $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
CVR $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ (through additional covariates) $\sqrt{}$
WET $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ (through weighted regression) $\sqrt{}$
Note: For PSW, survey weights are used to estimate the propensity score $e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$ and construct the balancing weights $\omega_{z,p}(\mathbf{X}_{i})=h(e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))/(p_{z}(\mathbf{X}_{i})e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))$; for MOM, survey weights are similarly used for $e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$ and $\omega_{z,p}(\mathbf{X}_{i})$, but the outcome regression does not include survey weights; for CVR, balancing weights $\omega_{z,p}(\mathbf{X}_{i})$ are included as additional covariates in the outcome regression; for WET, balancing weights $\omega_{z,p}(\mathbf{X}_{i})$ are incorporated as regression weights, introducing survey weights directly into the outcome regression process.
3.5 Sandwich variance estimation for augmented estimator under survey observational data
According to Proposition 3, all three augmented estimators are derived from the solution to the augmented weighting estimating equation 11. This allows us to construct a unified empirical sandwich variance estimator using M-estimation theory,[34] which enables computationally efficient solution for variance and confidence interval estimation under the context of survey observational data.
@@ 3.5 Sandwich variance estimation for augmented estimator under survey observational data
Specifically, the generalized form of the augmented estimator in Equation 11 can be expressed as $\widehat{\tau}_{h}^{\text{AW}}=\widehat{v}_{1}+\widehat{v}_{2}-\widehat{v}_{3}$. Adapting the empirical sandwich variance method [7, 9, 27] based on M-estimation theory, [29] we construct the estimating equation 19 for augmented estimators under the context of survey observational data. This estimating equation adjusts for the uncertainty in estimating the survey-weighted propensity score and two potential outcome models by utilizing score functions corresponding to $\boldsymbol{\theta}=\left(v_{1},v_{2},v_{3},\boldsymbol{\alpha}_{0}^{T},\boldsymbol{\alpha}_{1}^{T},\boldsymbol{\beta}_{\text{fp}}^{T},\boldsymbol{\beta}_{\text{sp}}^{T}\right)^{T}$, where $v_{1}$, $v_{2}$, and $v_{3}$ are the components in true value of the general augmented estimator $\tau_{h}^{\text{AW}}$, $\boldsymbol{\alpha}_{0}$ and $\boldsymbol{\alpha}_{1}$ are the parameters of the outcome models $m_{0}^{\text{AW}}(\mathbf{X};\boldsymbol{\alpha}_{0})$ and $m_{1}^{\text{AW}}(\mathbf{X};\boldsymbol{\alpha}_{1})$, respectively, and $\boldsymbol{\beta}_{\text{fp}}$ and $\boldsymbol{\beta}_{\text{sp}}$ represent the parameters of the sample-level and population-level propensity score models, respectively. Therefore, the estimating equation for the augmented estimators under the context of survey observational data is:
@@
$\mathbf{0}=\sum_{i=1}^{n}\boldsymbol{\Psi}_{\boldsymbol{\theta}}\left(\mathbf{X}_{i},Y_{i},Z_{i}\right)=\sum_{i=1}^{n}\left[\begin{array}[]{c}\psi_{\text{PS,FP}}\\ \psi_{\text{PS,SP}}\\ \psi_{\text{OR},1}\\ \psi_{\text{OR},0}\\ \psi_{\text{v1}}\\ \psi_{\text{v2}}\\ \psi_{\text{v3}}\end{array}\right]=\sum_{i=1}^{n}\left[\begin{array}[]{l}\omega_{\text{PS,FP}}\left(\mathbf{X}_{i}\right)\mathbf{X}_{i}\left(Z_{i}-e_{\text{fp}}\left(\mathbf{X}_{i};\boldsymbol{\beta}_{\text{fp}}\right)\right)\\ \omega_{\text{PS,SP}}\left(\mathbf{X}_{i}^{*}\right)\mathbf{X}_{i}^{*}\left(Z_{i}-e_{\text{sp}}\left(\mathbf{X}_{i}^{*};\boldsymbol{\beta}_{\text{sp}}\right)\right)\\ \omega_{\text{OR},1}\left(\mathbf{X}_{i}^{*}\right)Z_{i}\mathbf{X}_{i}^{\prime}\left(Y_{i}-m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)\right)\\ \omega_{\text{OR},0}\left(\mathbf{X}_{i}^{*}\right)(1-Z_{i})\mathbf{X}_{i}^{\prime\prime}\left(Y_{i}-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)\right)\\ h_{p}(\mathbf{X}_{i}^{*})\left(m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)-v_{1}\right)\\ \omega_{1,p}(\mathbf{X}_{i}^{*})Z_{i}\left(Y_{i}-m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)-v_{2}\right)\\ \omega_{0,p}(\mathbf{X}_{i}^{*})(1-Z_{i})\left(Y_{i}-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)-v_{3}\right)\end{array}\right]$ (19)
@@
The equation above includes six distinct components: $\psi_{\text{PS,FP}}$, $\psi_{\text{PS,SP}}$, $\psi_{\text{OR},1}$, $\psi_{\text{OR},0}$, $\psi_{\text{v1}}$, $\psi_{\text{v2}}$, and $\psi_{\text{v3}}$, each representing different aspects of the augmented weighting estimation process. The first two components, $\psi_{\text{PS,FP}}$ and $\psi_{\text{PS,SP}}$, correspond to the score functions associated with the estimation of the propensity score model at the sample and population levels, respectively. In retrospective settings, $e_{\text{fp}}(\mathbf{X})$ is used exclusively to calculate $r_{z}(\mathbf{X}_{i})$, which approximates $P(S_{i}=1\mid\mathbf{X}_{i})=r_{z}(\mathbf{X}_{i})p_{z}(\mathbf{X}_{i})$, and is applied in the augmented estimator to tilt the imputed outcome difference. The term $\omega_{\text{PS,FP}}(\mathbf{X}_{i})$ represents the regression weights for estimating the within-sample propensity score and is set to $1$. Furthermore, $\omega_{\text{PS,SP}}(\mathbf{X}_{i}^{*})$ refers to the regression weights applied to the covariate vector $\mathbf{X}_{i}^{*}$ in the population level propensity score model, which is either the survey weights when the regression is survey-weighted (W.PS) or set to $1$ when the regression is unweighted (U.PS). Depending on the model specification, $\mathbf{X}_{i}^{*}$ may consist of baseline covariates alone or specified original covariates with survey weights included as an additional covariate (C.PS), as recommended by DuGoff et al.[3], to capture unique design features and address latent sampling information , enhancing covariate balance and reducing bias from unmeasured confounders.
@@
The components $\psi_{\text{OR},1}$ and $\psi_{\text{OR},0}$ represent the score functions for estimating the unobserved potential outcome models for the treated ($Z_{i}=1$) and control ($Z_{i}=0$) groups, respectively. The weights $\omega_{\text{OR},1}(\mathbf{X}_{i}^{*})$ and $\omega_{\text{OR},0}(\mathbf{X}_{i}^{*})$ depend on the choice of augmented estimator. For the moment estimator (MOM) and clever covariate estimator (CVR), these weights are set to 1. For the weighted regression estimator (WET), they correspond to the balancing weights $\omega_{1,p}(\mathbf{X}_{i}^{*})$ and $\omega_{0,p}(\mathbf{X}_{i}^{*})$ as specified in Section 3.4. The outcome models $m_{1}^{\text{AW}}$ and $m_{0}^{\text{AW}}$ reflect the specific type of augmentation used, where the superscript “AW” can denote ”MOM”, ”CVR”, or ”WET”, depending on the choice of estimator. For the clever covariate estimator, the outcome model covariates $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$ include the clever covariates for each arm, whereas for MOM and WET, only the baseline covariates are included, i.e., $\mathbf{X}_{i}$.
@@
The component $\psi_{v_{1}}$ pertains to the difference in predicted full outcomes between the treated and control models, adjusted by the tilting function $h_{p}(\mathbf{X}_{i}^{*})=h(e_{\text{sp}}(\mathbf{X}_{i}^{*};\boldsymbol{\beta}))/P(S_{i}=1\mid\mathbf{X}_{i})$ and accounting for the first component $v_{1}$ in the general augmented treatment effect estimation 11. Here, $m_{1}^{\text{AW}}(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1})$ and $m_{0}^{\text{AW}}(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0})$ represent the specified outcome models for the treated and control groups, respectively, using covariate vectors $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$. The components $\psi_{v_{2}}$ and $\psi_{v_{3}}$ represent the residual adjustments for the treated and control groups, respectively. These adjustments incorporate the balancing weights $\omega_{1,p}(\mathbf{X}_{i}^{*})$ and $\omega_{0,p}(\mathbf{X}_{i}^{*})$, which are designed to adjust for confounding effects. These balancing weights are as described in Table 1 and are generalized for survey observational data under different model specifications.
@@
Table 3: Covariate vector notations used in the estimating equations for the sandwich variance estimation of augmented estimators under survey observational data.
@@
$\mathbf{X}_{i}^{*}$ Covariates used in the propensity score model $e_{\text{sp}}(\mathbf{X}_{i}^{*};\boldsymbol{\beta})$. In unweighted (U.PS) and survey-weighted regression scenarios (W.PS), it includes only baseline covariates. When survey weights are incorporated as an additional covariate (C.PS), it contains both the specified original covariates and the survey weights.
$\mathbf{X}_{i}^{\prime}$ Covariates in the treated outcome regression model $m_{1}^{\text{AW}}(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1})$. For CVR, this includes clever covariates; for MOM and WET, only baseline covariates are used.
$\mathbf{X}_{i}^{\prime\prime}$ Covariates in the control outcome regression model $m_{0}^{\text{AW}}(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0})$. Mirrors the definition of $\mathbf{X}_{i}^{\prime}$ based on the estimator type.
Note: The specified original covariates used in the propensity score model may differ from those in the outcome regression models, and the covariates used in the outcome models for each arm are not necessarily the same. In MOM and WET, where both outcome models include only baseline covariates without distinct clever covariates, $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$ may be identical or even unified, using one model for both arms’ counterfactual outcome predictions. In cases where the same set of covariates are used across all models, all three covariate notations ($\mathbf{X}_{i}^{*}$, $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$) in MOM and WET with U.PS or W.PS could be identical.

### v2

@@ 3.4 Weighted regression estimator
A third augmented estimator is the weighted regression estimator (WET)[28, 27, 29, 30] that, in our case, utilizes balancing weights during the outcome model fitting process. Adapted to our setting of survey sampling, the weighted regression estimator can be expressed as,
@@
$\widehat{\tau}_{h}^{\text{WET}}=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(\mathbf{X}_{i})\left(m_{1}^{\text{WET}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{WET}}\right)-m_{0}^{\text{WET}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{WET}}\right)\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})},$ (19)
where $m_{1}^{\text{WET}}\left(\mathbf{X};\boldsymbol{\alpha}_{1}^{\text{WET}}\right)$ and $m_{0}^{\text{WET}}\left(\mathbf{X};\boldsymbol{\alpha}_{0}^{\text{WET}}\right)$ are the outcome regression models with canonical link functions, similar to those in the clever covariate estimator, but using weighted regression with balancing weights $\omega_{1,p}\left(\mathbf{X}\right)$ and $\omega_{0,p}\left(\mathbf{X}\right)$. In this estimator, survey weights must be incorporated in three places: estimating the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$, constructing the balancing weights $\omega_{z,p}(\mathbf{X}_{i})$ that serve as regression weights in the outcome modeling, and the ratio $h(\mathbf{X}_{i})/p(\mathbf{X}_{i})$ for standardizing predictions to the target population. The weighted regression estimator is asymptotically equivalent to the one-step moment and clever covariate estimators, sharing the same influence function and robustness properties. When $h(\mathbf{X})\propto 1$, the estimator achieves double robustness. When $h(\mathbf{X})$ depends on $e_{\text{sp}}(\mathbf{X})$, the estimator is singly robust with respect to the propensity score model; if the propensity score model is misspecified, the estimator converges to the shifted estimand $\widetilde{\tau}_{h}$ as discussed in Section 3.1.
@@
Compared to the clever covariate estimator, the weighted regression estimator may offer better performance in scenarios involving highly variable survey weights or extreme propensity scores. As discussed by Kang et al.[26] and Robins et al.[27], the weighted regression estimator is more stable under extreme weights. However, under the IPW scheme, the weighted regression estimator remains susceptible to the impact of extreme inverse propensity scores, which can lead to inflated variance and finite-sample bias, particularly when these extreme weights dominate the regression model. We plan to examine to what extent these observations carry forward to the survey observational study settings in Section 4.
For ease of reference, Table 2 summarizes where survey weights need to be applied for each simple weighting estimator and augmented weighting estimator.
@@
Estimator Estimating propensity scores Fitting the outcome regression model Standardizing outcomes to the target estimand Standardizing weighted residuals
PSW $\sqrt{}$ $\times$ $\sqrt{}$ $\times$
MOM $\sqrt{}$ $\times$ $\sqrt{}$ $\sqrt{}$
CVR $\sqrt{}$ $\sqrt{}$ (clever covariates) $\sqrt{}$ $\times$
WET $\sqrt{}$ $\sqrt{}$ (weighted regression) $\sqrt{}$ $\times$
Note: All estimators use survey-weighted logistic regression to estimate the population-level propensity score $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$. PSW uses balancing weights $\omega_{z,p}(\mathbf{X}_{i})=h(\mathbf{X}_{i})/(p_{z}(\mathbf{X}_{i})e_{\text{sp}}(\mathbf{X}_{i}))$ to standardize observed outcomes for each group, while augmented estimators use the ratio $h(\mathbf{X}_{i})/p(\mathbf{X}_{i})$ to standardize predicted full potential outcomes, achieving the target estimand. Additionally, CVR incorporates balancing weights $\omega_{z,p}(\mathbf{X}_{i})$ as clever covariates in the outcome model, WET uses $\omega_{z,p}(\mathbf{X}_{i})$ as regression weights, and MOM fits an unweighted outcome model but uses $\omega_{z,p}(\mathbf{X}_{i})$ to standardize its weighted residuals to achieve additional robustness.
3.5 Variance estimation for augmented estimators
@@
The generalized form of the augmented estimator in Equation (13) can be expressed as $\widehat{\tau}_{h}^{\text{AW}}=\widehat{v}_{1}+\widehat{v}_{2}-\widehat{v}_{3}$. Leveraging the M-estimation theory, we construct the joint estimating equations (20) for sandwich variance estimation. These estimating equations adjust for the uncertainty in estimating the propensity score and outcome models by utilizing score functions corresponding to $\boldsymbol{\theta}=\left(v_{1},v_{2},v_{3},\boldsymbol{\alpha}_{0}^{T},\boldsymbol{\alpha}_{1}^{T},\boldsymbol{\beta}_{\text{fp}}^{T},\boldsymbol{\beta}_{\text{sp}}^{T}\right)^{T}$, where $v_{1}$, $v_{2}$, and $v_{3}$ are the components of the general augmented estimator $\tau_{h}^{\text{AW}}$, $\boldsymbol{\alpha}_{0}$ and $\boldsymbol{\alpha}_{1}$ are the parameters of the outcome models $m_{0}^{\text{AW}}(\mathbf{X};\boldsymbol{\alpha}_{0})$ and $m_{1}^{\text{AW}}(\mathbf{X};\boldsymbol{\alpha}_{1})$, respectively, and $\boldsymbol{\beta}_{\text{fp}}$ and $\boldsymbol{\beta}_{\text{sp}}$ represent the parameters of the sample-level and population-level propensity score models, respectively.
@@
$\mathbf{0}=\sum_{i=1}^{n}\boldsymbol{\Psi}_{\boldsymbol{\theta}}\left(\mathbf{X}_{i},Y_{i},Z_{i}\right)=\sum_{i=1}^{n}\left[\begin{array}[]{c}\psi_{\text{PS,FP}}\\ \psi_{\text{PS,SP}}\\ \psi_{\text{OR},1}\\ \psi_{\text{OR},0}\\ \psi_{\text{v1}}\\ \psi_{\text{v2}}\\ \psi_{\text{v3}}\end{array}\right]=\sum_{i=1}^{n}\left[\begin{array}[]{l}\omega_{\text{PS,FP}}\left(\mathbf{X}_{i}\right)\mathbf{X}_{i}\left(Z_{i}-e_{\text{fp}}\left(\mathbf{X}_{i};\boldsymbol{\beta}_{\text{fp}}\right)\right)\\ \omega_{\text{PS,SP}}\left(\mathbf{X}_{i}^{*}\right)\mathbf{X}_{i}^{*}\left(Z_{i}-e_{\text{sp}}\left(\mathbf{X}_{i}^{*};\boldsymbol{\beta}_{\text{sp}}\right)\right)\\ \omega_{\text{OR},1}\left(\mathbf{X}_{i}^{*}\right)Z_{i}\mathbf{X}_{i}^{\prime}\left(Y_{i}-m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)\right)\\ \omega_{\text{OR},0}\left(\mathbf{X}_{i}^{*}\right)(1-Z_{i})\mathbf{X}_{i}^{\prime\prime}\left(Y_{i}-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)\right)\\ h(\mathbf{X}_{i}^{*})/p(\mathbf{X}_{i})\left(m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)-v_{1}\right)\\ \omega_{1,p}(\mathbf{X}_{i}^{*})Z_{i}\left(Y_{i}-m_{1}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1}\right)-v_{2}\right)\\ \omega_{0,p}(\mathbf{X}_{i}^{*})(1-Z_{i})\left(Y_{i}-m_{0}^{\text{AW}}\left(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0}\right)-v_{3}\right)\end{array}\right].$ (20)
@@
The equation above includes seven distinct components: $\psi_{\text{PS,FP}}$, $\psi_{\text{PS,SP}}$, $\psi_{\text{OR},1}$, $\psi_{\text{OR},0}$, $\psi_{\text{v1}}$, $\psi_{\text{v2}}$, and $\psi_{\text{v3}}$, each representing different aspects of the augmented weighting estimation process. The first two components, $\psi_{\text{PS,FP}}$ and $\psi_{\text{PS,SP}}$, correspond to the score functions associated with the estimation of the propensity score model at the sample and population levels, respectively. In retrospective settings, $e_{\text{fp}}(\mathbf{X};\boldsymbol{\beta}_{\text{fp}})$ is used exclusively to calculate $r_{z}(\mathbf{X}_{i})$ for obtaining the marginalized sampling probability $p(\mathbf{X}_{i})=r_{z}(\mathbf{X}_{i})p_{z}(\mathbf{X}_{i})$, which is used for tilting the imputed outcome difference in the augmented estimator as discussed in Section 3.1. The term $\omega_{\text{PS,FP}}(\mathbf{X}_{i})$ represents the regression weights for estimating the within-sample propensity score and is set to $1$. Furthermore, $\omega_{\text{PS,SP}}(\mathbf{X}_{i}^{*})$ refers to the regression weights applied to the covariate vector $\mathbf{X}_{i}^{*}$ in the population-level propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta}_{\text{sp}})$, which is either the survey weights when the regression is survey-weighted (W.PS) or set to $1$ when the regression is unweighted (U.PS). The covariate vector $\mathbf{X}_{i}^{*}$ may consist of baseline covariates alone or may additionaly include survey weights as a covariate (C.PS), as recommended by DuGoff et al. [3] to capture unique design features and address latent sampling information so as to enhance the covariate balance.
@@
The components $\psi_{\text{OR},1}$ and $\psi_{\text{OR},0}$ represent the score functions for the outcome models for the treated ($z_{i}=1$) and control ($z_{i}=0$) groups, respectively. The regression weights $\omega_{\text{OR},1}(\mathbf{X}_{i}^{*})$ and $\omega_{\text{OR},0}(\mathbf{X}_{i}^{*})$ depend on the choice of augmented estimator: for the one-step moment estimator (MOM) and clever covariate estimator (CVR), these weights are set to 1; for the weighted regression estimator (WET), they correspond to the balancing weights $\omega_{1,p}(\mathbf{X}_{i}^{*})$ and $\omega_{0,p}(\mathbf{X}_{i}^{*})$ as specified in Section 3.4. The outcome models $m_{1}^{\text{AW}}$ and $m_{0}^{\text{AW}}$ reflect the specific type of augmentation used, where the superscript “AW” can denote ”MOM”, ”CVR”, or ”WET”. For the clever covariate estimator, the outcome model covariates $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$ include the clever covariates for each arm, whereas for MOM and WET, only the baseline covariates are used, i.e., $\mathbf{X}_{i}$.
@@
The component $\psi_{v_{1}}$ corresponds to the first component $v_{1}$ in the general augmented treatment effect estimator (13), representing the difference in predicted potential outcomes between the treated and control groups, adjusted by the ratio $h(\mathbf{X}_{i}^{*})/p(\mathbf{X}_{i})$, where the tilting function $h(\mathbf{X}_{i}^{*})$ may depend on $e_{\text{sp}}(\mathbf{X}_{i}^{*};\boldsymbol{\beta}_{\text{sp}})$ as specified in Table 1. The components $\psi_{v_{2}}$ and $\psi_{v_{3}}$ represent the residual adjustments for the treated and control groups, respectively, incorporating the balancing weights $\omega_{1,p}(\mathbf{X}_{i}^{*})$ and $\omega_{0,p}(\mathbf{X}_{i}^{*})$ to adjust for confounding, as described in Table 1.
@@
Table 3: Covariate vector notations used in the estimating equations for the sandwich variance estimation of augmented estimators for survey observational data.
@@
$\mathbf{X}_{i}^{*}$ Covariates used in the propensity score model $e_{\text{sp}}(\mathbf{X}_{i}^{*};\boldsymbol{\beta}_{\text{sp}})$. For unweighted (U.PS) and survey-weighted regression scenarios (W.PS), it includes only baseline covariates. When survey weights are incorporated as an additional covariate (C.PS), it contains both the baseline covariates and the survey weights.
$\mathbf{X}_{i}^{\prime}$ Covariates in the treated outcome regression model $m_{1}^{\text{AW}}(\mathbf{X}_{i}^{\prime};\boldsymbol{\alpha}_{1})$. For CVR, it includes clever covariates; for MOM and WET, only baseline covariates are used.
$\mathbf{X}_{i}^{\prime\prime}$ Covariates in the control outcome regression model $m_{0}^{\text{AW}}(\mathbf{X}_{i}^{\prime\prime};\boldsymbol{\alpha}_{0})$. Defined analogously to $\mathbf{X}_{i}^{\prime}$ based on the estimator type.
Note: The specified original covariates used in the propensity score model may differ from those in the outcome regression models, and the covariates used in the outcome models for each arm are not necessarily the same. In MOM and WET, where both outcome models include only baseline covariates without distinct clever covariates, $\mathbf{X}_{i}^{\prime}$ and $\mathbf{X}_{i}^{\prime\prime}$ may be identical or even unified, using one model for both arms’ counterfactual outcome predictions. In cases where the same covariates are used across all models, such as in MOM and WET with U.PS or W.PS, all three covariate notations ($\mathbf{X}_{i}^{*}$, $\mathbf{X}_{i}^{\prime}$, and $\mathbf{X}_{i}^{\prime\prime}$) could be identical.